# Generalized Guarantees for Variational Inference in the Presence of Even and Elliptical Symmetry

### v1

@@ 3.2 Guarantees from elliptical symmetry
Next we generalize earlier guarantees for VI with elliptical symmetries (Margossian and Saul,, 2025) to the broader family of $\varphi$-divergences. To begin, we note that if $p$ is elliptically symmetric, then it admits a point of even symmetry, $\mu$, and a normalized covariance matrix, $M$. In this case, there exists a spherically symmetric density $p_{0}$ satisfying
@@
Let $\mathcal{Q}$ be a location-scale family, and let $D_{\varphi}$ be a $\varphi$-divergence with strictly decreasing $\varphi$. Suppose $p$ is somewhere-strictly log concave over $\mathbb{R}^{d}$ and elliptically symmetric about $\mu$ with normalized covariance matrix $M$; also, suppose $f$ (as defined above) is everywhere continuously differentiable with $|f^{\prime}(0)|\!<\!\infty$. Then $D_{\varphi}(p||q)$ has a unique minimizer with respect to the location-scale parameters $(\nu,S)$ of $\mathcal{Q}$ at $\nu\!=\!\mu$ and $S\!=\!\gamma^{2}M$ for some $\gamma\in\mathbb{R}$.
A full proof is given in Appendix B.2. The main idea is to rewrite $D_{\varphi}(p||q_{\nu,S})$ as an integral over spherically symmetric functions in the transformed variable $\zeta\!=\!S^{-\frac{1}{2}}(z\!-\!\nu)$. Then, with some technical machinery, it can be shown that a unique stationary point of $D_{\varphi}(p||q_{\nu,S})$ occurs at $\nu\!=\!\mu$ and $S\!=\!\gamma^{2}M$, thus recovering the exact mean and correlation matrix.
As before, the theorem allows for misspecifications in $\mathcal{Q}$. Fig. 2 (right) illustrates VI’s ability to recover these statistics when $p$ is a student-t distribution and $\mathcal{Q}$ is the family of multivariate Gaussians. Though each divergence in Table 1 yields a different approximation, they all recover the mean and correlation of $p$. The theorem predicts this result for the reverse KL divergence and the Renyi-divergence with $\alpha\!\in\!(0,1)$.
@@ 3.3 Guarantees from partial symmetry
Let $\mathcal{Q}$ be a location family. If $p$ is even symmetric along $\sigma$ with a constant point of even symmetry, then $\nu_{\sigma}\!=\!\mu_{\sigma}$ is a stationary point of $D_{\varphi}(p||q_{\nu})$ for any fixed $\nu_{\bar{\sigma}}$. Also, if $\varphi$ is strictly decreasing and $p(z_{\sigma}|z_{\bar{\sigma}})$ is somewhere-strictly log concave over $\mathbb{R}^{|\sigma|}$, then $\nu_{\sigma}\!=\!\mu_{\sigma}$ at all stationary points of $D_{\varphi}(p||q_{\nu})$.
A complete proof is given in Appendix B.3, and its steps closely follow those for the proof of Theorem 10.
We now provide guarantees for VI when $p$ is elliptically symmetric along $\sigma$. For these guarantees, we must make a number of additional assumptions—namely, that $\mathcal{Q}$ is the family of multivariate Gaussians, that the variational approximation is found by minimizing $\text{KL}(q||p)$ (as opposed to any $\varphi$-divergence), and also that the conditional mean and normalized covariance of $p(z_{\sigma}|z_{\bar{\sigma}})$ do not depend on $z_{\bar{\sigma}}$. (The utility of these assumptions is made clear in the proof of the next theorem.) Despite these restrictions, the theorem still has a fairly broad scope: it covers the most common setting of black box VI (e.g. Kucukelbir et al.,, 2017), where $\text{KL}(q||p)$ is minimized to find the best multivariate Gaussian approximation $q$, and it also covers non-trivial target densities such as the Rosenbrock distribution (Roberts et al.,, 1997) and the funnel of eq. (6) and some extensions we explore in Appendix B.5. The latter often characterizes hierarchical priors in Bayesian models.
@@
Let $\mathcal{Q}$ be the family of multivariate Gaussians. Suppose $p$ is elliptically symmetric along $\sigma$ about a constant point of even symmetry and with a constant normalized covariance matrix, and suppose $f$ is continuously differentiable and $|f^{\prime}(0)|\!<\!\infty$. Also suppose $p(z_{\sigma}|z_{\bar{\sigma}})$ is somewhere-strictly log concave over $\mathbb{R}^{|\sigma|}$. Then any minimizer of $\text{KL}(q||p)$ satisfies $\nu_{\sigma}\!=\!\mu_{\sigma}$ and $S_{\sigma\sigma}\!=\!\gamma^{2}\Sigma_{\sigma\sigma}$ for some unique $\gamma\in\mathbb{R}$.
@@
Note that the reverse KL divergence, for which $\varphi(t)$ is linear in Table 1, is the only strictly decreasing $\varphi$-divergence that permits such a decomposition. The keystone of the proof is to find values for $(\nu_{\sigma},S_{\sigma\sigma},S_{\sigma\bar{\sigma}})$ that uniformly minimize the integrand of eq. (14) for all values of $z_{\bar{\sigma}}$, and then to note that this minimum is realized for any values of $(\nu_{\bar{\sigma}},S_{\bar{\sigma}\bar{\sigma}})$ that fix $q(z_{\bar{\sigma}})$ and hence $\text{KL}(q_{\bar{\sigma}}\,||\,p_{\bar{\sigma}})$ in the remaining term.
@@
Table 2: Target distributions for experiments and estimated asymmetry $\alpha_{\text{90}}$; see eq. (19). The hierarchical models (schools, disease, SKIM) were implementing using (i) a standard centered parameterization, (ii) a non-centered parameterization with less asymmetry, (iii) an approximate marginalization of the latent variables, yielding a collapsed (and even less asymmetric) posterior; see the discussion in section 4.1.
@@
Since $p$ has a constant point of even symmetry, the right side of eq. (15) cannot depend on $z_{\bar{\sigma}}$, and so we set $S_{\sigma\bar{\sigma}}={\bf 0}$. Then eq. (15) is solved by $\nu_{\sigma}\!=\!{\rm E}_{p}[z_{\sigma}|z_{\bar{\sigma}}]$ and eq. (16) by $S_{\sigma\sigma}\!\propto\!{\rm Cov}_{p}[z_{\sigma}|z_{\bar{\sigma}}]$. Finally, we show in Lemma 18 of Appendix B.4 that $\mu_{\sigma}\!=\!{\rm E}_{p}[z_{\sigma}|z_{\bar{\sigma}}]$ and $\Sigma_{\sigma\sigma}\!\propto\!{\rm Cov}_{p}[z_{\sigma}|z_{\bar{\sigma}}]$ if $p$ is elliptically symmetric along $\sigma$ with a constant point of even symmetry and a constant normalized covariance matrix. This completes the proof. ∎
Some of the conditions for this proof can be relaxed. In Appendix B.4, we consider the case where $\mathbb{E}_{p}[z_{\sigma}|z_{\bar{\sigma}}]$ is a linear function of $z_{\bar{\sigma}}$ and show that Gaussian VI provably recovers the conditional mean and correlations. In Appendix B.5, we apply these theoretical results to the funnel of eq. (6) and variations thereof.
@@ 4 NUMERICAL EXPERIMENTS
We investigate the performance of VI on a diverse set of target densities (Table 2). The first three are synthetic—a multivariate student-t, an elliptical funnel, and a crescent distribution—and the others are derived from Bayesian hierarchical models, including a model of education data (Gelman et al.,, 2013), a sparse kernel interaction model of gene microarray data (Agrawal et al.,, 2019), and a Gaussian process model of mortality counts with a Poisson likelihood (Vanhatalo et al.,, 2010). The full definition of each model is provided in Appendix C.1.
@@ 4.1 Are Bayesian posteriors symmetric?
We now discuss how approximate symmetries may arise in a Bayesian analysis. Given a latent variable $z$ and observation $x$, the posterior is given by $\pi(z|x)\!\propto\!\pi(z)\,\pi(x|z)$, and the likelihood often factors as $\pi(x|z)\!=\!\prod_{i}\pi(x_{i}|z)$. In general this product may not have an even or elliptical symmetry even when one is found in each of its terms. However, if the product is dominated by a symmetric prior (for sparse data) or the likelihood (for rich data), then the posterior tends to be approximately symmetric (Margossian and Saul,, 2025). In the latter regime, the Bernstein-von Mises theorem (van der Vaart,, 1998) is often invoked to argue the posterior is approximately Gaussian.
@@
as in Gaussian processes and, for example, the disease and SKIM targets in Table 2. But there are also strategies to mitigate these sources of asymmetry, without changing the generative model. One strategy is to use a non-centered parameterization (Papaspiliopoulos et al.,, 2007). Let $L$ denote the Cholesky decomposition of the prior covariance matrix in eq. (17), such that $LL^{T}\!=\!\exp(2\tau)C(\rho)$. This strategy introduces an auxiliary variable $\varepsilon\!\sim\!\mathcal{N}(0,1)$ and recomputes the likelihood (equivalently) as $\pi(x|\varepsilon)=\pi(x|\theta\!=\!L\varepsilon+\mu)$. A Bayesian inference algorithm then approximates the posterior $\pi(\varepsilon,\mu,\tau,\rho|x)$, where the funnels in the prior have been removed due to the independence of $\varepsilon$ and $\tau$. A caveat is that this non-centered parameterization can complicate the likelihood, especially in rich data regimes, leading to a challenging posterior geometry.
@@
Here the marginalization serves to remove the funnel over $\theta$ and $\tau$ in the prior. Inference on $\theta$ is performed post-hoc by approximating $\pi(\theta|\mu,\tau,\rho,x)$. If the likelihood is normal, then the marginalization is tractable; otherwise it must be approximated—for example, via an integrated Laplace approximation (e.g., Rasmussen and Williams,, 2006; Rue et al.,, 2009).
@@ 4.2 VI algorithm
We use automatic differentiation VI (ADVI, Kucukelbir et al.,, 2017), as implemented in Stan (Carpenter et al.,, 2017; Stan Development Team,, 2025). ADVI employs a Gaussian approximation and minimizes $\text{KL}(q||p)$ via stochastic optimization. Details on the implementation of ADVI are described in Appendix C.2. While our theorems consider more general $\varphi$-divergences, we do not have a reliable way to minimize these divergences, and so we only give results for $\text{KL}(q||p)$. For models with a collapsed posterior, we use Stan’s prototype integrated Laplace approximation (Margossian et al.,, 2020). As a benchmark, we approximate $\pi(z|x)$ with exact sampling of the synthetic targets and long runs (20000 iterations) of Stan’s Hamiltonian Monte Carlo sampler (Hoffman and Gelman,, 2014; Betancourt,, 2018) with a non-centered parameterization for the hierarchical models.
@@ 4.3 Results
We find a similar trend for estimates of correlations; generally VI yields better estimates in models with more symmetry. This trend is clear for the synthetic targets, less so for the hierarchical models, perhaps because the latter have many nearly zero correlations. We provide these results in Appendix C.3.
@@ 5 Discussion
In this paper we derive novel symmetry-based guarantees for VI with a broad class of divergences and in cases where $p$ exhibits partial symmetries. Our results provide not only theoretical insight, but also prescriptions for practitioners using VI to approximate the posteriors of Bayesian hierarchical models. They suggest, in particular, that VI can be improved by implementing these models in certain ways. These prescriptions are reminiscent of those known to improve the performance of MCMC samplers on challenging posteriors (e.g., Betancourt and Girolami,, 2015; Gómez-Rubio and Rue,, 2018; Margossian et al.,, 2020)
The above considerations suggest one way to improve VI in Bayesian hierarchical models. But a more common approach is simply to choose a richer variational family $\mathcal{Q}$—for example, one allowing skewed approximations (Tan and Chen,, 2024), or even semi-parametric or non-parametric approximations (e.g., Agrawal et al.,, 2020; Xu et al.,, 2023; Cai et al., 2024a, ; Xu and Campbell,, 2025). While this approach requires less bespoke manipulations of $p$, its computational expense grows quickly with the complexity of $\mathcal{Q}$. It is therefore of interest to understand how well simpler variational families can perform. Moving forward, we advocate a workflow in which VI proceeds first with a restricted family $\mathcal{Q}$ (e.g., factorized or location-scale), the accuracy of the inference is checked (here, it would be interesting to develop a symmetry-based check), and then slight corrections are applied (e.g., Yao et al.,, 2018; Giordano et al.,, 2018) and $\mathcal{Q}$ is progressively complexified as necessary.

### v2

@@ 3.2 Guarantees from elliptical symmetry
Next, we present guarantees for VI with elliptical symmetries. Our results are stated in terms of $f$-divergences, however the conditions we impose on $\varphi$, and hence on $f$, restrict the application of our results to $\text{KL}(q||p)$ in Table 1, and so our results only provide a nominal generalization of the guarantees derived by Margossian:2025robust. Still, we present the results for completeness.
To begin, we note that if $p$ is elliptically symmetric, then it admits a point of even symmetry, $\mu$, and a normalized covariance matrix, $M$. In this case, there exists a spherically symmetric density $p_{0}$ satisfying
@@
Let $\mathcal{Q}$ be a location-scale family, and let $D_{f}$ be an $f$-divergence with $\varphi(v)=f\circ\exp(v)$ convex and strictly decreasing. Suppose $p$ is somewhere-strictly log concave over $\mathbb{R}^{d}$ and elliptically symmetric about $\mu$ with normalized covariance matrix $M$; also, suppose $g$ (as defined above) is everywhere continuously differentiable with $|g^{\prime}(0)|\!<\!\infty$. Then $D_{f}(p||q)$ has a unique minimizer with respect to the location-scale parameters $(\nu,S)$ of $\mathcal{Q}$ at $\nu\!=\!\mu$ and $S\!=\!\gamma^{2}M$ for some $\gamma\in\mathbb{R}$.
A full proof is given in Appendix A.2. The main idea is to rewrite $D_{\varphi}(p||q_{\nu,S})$ as an integral over spherically symmetric functions in the transformed variable $\zeta\!=\!S^{-\frac{1}{2}}(z\!-\!\nu)$. Then, with some technical machinery, it can be shown that a unique stationary point of $D_{\varphi}(p||q_{\nu,S})$ occurs at $\nu\!=\!\mu$ and $S\!=\!\gamma^{2}M$, thus recovering the exact mean and correlation matrix.
Fig. 2 (right) illustrates VI’s ability to recover these statistics when $p$ is a student-t distribution and $\mathcal{Q}$ is the family of multivariate Gaussians. Though each divergence in Table 1 yields a different approximation, they all recover the mean and correlation of $p$. As before, this experiment provides a stronger result than the theorem predicts, since exact recovery of the mean and correlation is achieved by all divergences, and the target $p$ is quasi-concave, rather than log concave.
@@ 3.3 Guarantees from partial symmetry
Let $\mathcal{Q}$ be a location family. If $p$ is even symmetric along $\sigma$ with a constant point of even symmetry, then $\nu_{\sigma}\!=\!\mu_{\sigma}$ is a stationary point of $D_{f}(p||q_{\nu})$ for any fixed $\nu_{\bar{\sigma}}$. Also, if $\varphi(v)=f\circ\exp(v)$ is strictly decreasing and $p(z_{\sigma}|z_{\bar{\sigma}})$ is somewhere-strictly log concave over $\mathbb{R}^{|\sigma|}$, then $\nu_{\sigma}\!=\!\mu_{\sigma}$ at all stationary points of $D_{\varphi}(p||q_{\nu})$.
A complete proof is given in Appendix A.3, and its steps closely follow those for the proof of Theorem 10.
We now provide guarantees for VI when $p$ is elliptically symmetric along $\sigma$. For these guarantees, we must make a number of additional assumptions. First, we require $\varphi$ to be linear, an assumption which is only verified by $\text{KL}(q||p)$ in Table 1. Second, we take $\mathcal{Q}$ to be the family of mulitvariate Gaussians (as opposed to any location-scale family). Third, we assume that the conditional mean and normalized covariance matrix of $p(z_{\sigma}|z_{\bar{\sigma}})$ do not depend on $z_{\bar{\sigma}}$. (The utility of these assumptions is made clear in the proof of the next theorem.) Despite these restrictions, the theorem still has a fairly broad scope: it covers the most common setting of black box VI (e.g. Kucukelbir:2017), where $\text{KL}(q||p)$ is minimized to find the best multivariate Gaussian approximation $q$, and it also covers non-trivial target densities such as the Rosenbrock distribution (Roberts:1997) and the funnel of eq. (8) and some extensions we explore in Appendix A.5. The latter often characterizes hierarchical priors in Bayesian models.
@@
Let $\mathcal{Q}$ be the family of multivariate Gaussians. Suppose $p$ is elliptically symmetric along $\sigma$ about a constant point of even symmetry and with a constant normalized covariance matrix, and suppose $g$ is continuously differentiable and $|g^{\prime}(0)|\!<\!\infty$. Also suppose $p(z_{\sigma}|z_{\bar{\sigma}})$ is somewhere-strictly log concave over $\mathbb{R}^{|\sigma|}$. Then any minimizer of $\text{KL}(q||p)$ satisfies $\nu_{\sigma}\!=\!\mu_{\sigma}$ and $S_{\sigma\sigma}\!=\!\gamma^{2}\Sigma_{\sigma\sigma}$ for some unique $\gamma\in\mathbb{R}$.
@@
Note that the reverse KL divergence, for which $\varphi(u)$ is linear in Table 1, is the only $f$-divergence that permits such a decomposition. The keystone of the proof is to find values for $(\nu_{\sigma},S_{\sigma\sigma},S_{\sigma\bar{\sigma}})$ that uniformly minimize the integrand of eq. (16) for all values of $z_{\bar{\sigma}}$, and then to note that this minimum is realized for any values of $(\nu_{\bar{\sigma}},S_{\bar{\sigma}\bar{\sigma}})$ that fix $q(z_{\bar{\sigma}})$ and hence $\text{KL}(q_{\bar{\sigma}}\,||\,p_{\bar{\sigma}})$ in the remaining term.
@@
Table 2: Target distributions for experiments and estimated asymmetry $\alpha_{\text{90}}$; see eq. (21). The hierarchical models (schools, disease, SKIM) were implemented using (i) a standard centered parameterization, (ii) a non-centered parameterization with less asymmetry, (iii) an approximate marginalization of the latent variables, yielding a collapsed (and even less asymmetric) posterior; see the discussion in section 4.1.
@@
Since $p$ has a constant point of even symmetry, the right side of eq. (17) cannot depend on $z_{\bar{\sigma}}$, and so we set $S_{\sigma\bar{\sigma}}={\bf 0}$. Then eq. (17) is solved by $\nu_{\sigma}\!=\!{\rm E}_{p}[z_{\sigma}|z_{\bar{\sigma}}]$ and eq. (18) by $S_{\sigma\sigma}\!\propto\!{\rm Cov}_{p}[z_{\sigma}|z_{\bar{\sigma}}]$. Finally, we show in Lemma 17 of Appendix A.4 that $\mu_{\sigma}\!=\!{\rm E}_{p}[z_{\sigma}|z_{\bar{\sigma}}]$ and $\Sigma_{\sigma\sigma}\!\propto\!{\rm Cov}_{p}[z_{\sigma}|z_{\bar{\sigma}}]$ if $p$ is elliptically symmetric along $\sigma$ with a constant point of even symmetry and a constant normalized covariance matrix. This completes the proof. ∎
Some of the conditions for this proof can be relaxed. In Appendix A.4, we consider the case where $\mathbb{E}_{p}[z_{\sigma}|z_{\bar{\sigma}}]$ is a linear function of $z_{\bar{\sigma}}$ and show that Gaussian VI provably recovers the conditional mean and correlations. In Appendix A.5, we apply these theoretical results to the funnel of eq. (8) and variations thereof.
@@ 4 NUMERICAL EXPERIMENTS
We investigate the performance of VI on a diverse set of target densities (Table 2). The first three are synthetic—a multivariate student-t, an elliptical funnel, and a crescent distribution—and the others are derived from Bayesian hierarchical models, including a model of education data (Gelman:2013), a sparse kernel interaction model of gene microarray data (Agrawal:2019), and a Gaussian process model of mortality counts with a Poisson likelihood (Vanhatalo:2019). The full definition of each model is provided in Appendix B.1.
@@ 4.1 Are Bayesian posteriors symmetric?
We now discuss how approximate symmetries may arise in a Bayesian analysis. Given a latent variable $z$ and observation $x$, the posterior is given by $\pi(z|x)\!\propto\!\pi(z)\,\pi(x|z)$, and the likelihood often factors as $\pi(x|z)\!=\!\prod_{i}\pi(x_{i}|z)$. In general this product may not have an even or elliptical symmetry even when one is found in each of its terms. However, if the product is dominated by a symmetric prior (for sparse data) or the likelihood (for rich data), then the posterior tends to be approximately symmetric (Margossian:2025robust). In the latter regime, the Bernstein-von Mises theorem (Vaart:1998) is often invoked to argue the posterior is approximately Gaussian.
@@
as in Gaussian processes and, for example, the disease and SKIM targets in Table 2. But there are also strategies to mitigate these sources of asymmetry, without changing the generative model. One strategy is to use a non-centered parameterization (Papaspiliopoulos:2007). Let $L$ denote the Cholesky decomposition of the prior covariance matrix in eq. (19), such that $LL^{T}\!=\!\exp(2\tau)C(\rho)$. This strategy introduces an auxiliary variable $\varepsilon\!\sim\!\mathcal{N}(0,1)$ and recomputes the likelihood (equivalently) as $\pi(x|\varepsilon)=\pi(x|\theta\!=\!L\varepsilon+\mu)$. A Bayesian inference algorithm then approximates the posterior $\pi(\varepsilon,\mu,\tau,\rho|x)$, where the funnels in the prior have been removed due to the independence of $\varepsilon$ and $\tau$. A caveat is that this non-centered parameterization can complicate the likelihood, especially in rich data regimes, leading to a challenging posterior geometry.
@@
Here the marginalization serves to remove the funnel over $\theta$ and $\tau$ in the prior. Inference on $\theta$ is performed post-hoc by approximating $\pi(\theta|\mu,\tau,\rho,x)$. If the likelihood is normal, then the marginalization is tractable; otherwise it must be approximated—for example, via an integrated Laplace approximation (e.g., Rasmussen:2006; Rue:2009).
@@ 4.2 VI algorithm
We use automatic differentiation VI (ADVI, Kucukelbir:2017), as implemented in Stan (Carpenter:2017; Stan:2025). ADVI employs a Gaussian approximation and minimizes $\text{KL}(q||p)$ via stochastic optimization. Details on the implementation of ADVI are described in Appendix B.2. While our theorems consider more general $f$-divergences, we do not have a reliable way to minimize these divergences, and so we only give results for $\text{KL}(q||p)$. For models with a collapsed posterior, we use Stan’s prototype integrated Laplace approximation (Margossian:2020). As a benchmark, we approximate $\pi(z|x)$ with exact sampling of the synthetic targets and long runs (20000 iterations) of Stan’s Hamiltonian Monte Carlo sampler (Hoffman:2014; Betancourt:2018) with a non-centered parameterization for the hierarchical models.
@@ 4.3 Results
We find a similar trend for estimates of correlations; generally VI yields better estimates in models with more symmetry. This trend is clear for the synthetic targets, less so for the hierarchical models, perhaps because the latter have many nearly zero correlations. We provide these results in Appendix B.3.
@@ 5 Discussion
In this paper we provide some extension of symmetry-based guarantees for VI when minimizing $f$-divergences, although our stronger results still only apply to $\text{KL}(q||p)$. We also provide guarantees in cases where $p$ exhibits partial symmetries. Our results provide not only theoretical insight, but also prescriptions for practitioners using VI to approximate the posteriors of Bayesian hierarchical models. They suggest, in particular, that VI can be improved by implementing these models in certain ways. These prescriptions are reminiscent of those known to improve the performance of MCMC samplers on challenging posteriors (e.g., Betancourt:2015; Gomez:2018; Margossian:2020)
The above considerations suggest one way to improve VI in Bayesian hierarchical models. But a more common approach is simply to choose a richer variational family $\mathcal{Q}$—for example, one allowing skewed approximations (Tan:2024), or even semi-parametric or non-parametric approximations (e.g., Agrawal:2020; Xu:2023; Cai:2024b; Xu:2025). While this approach requires less bespoke manipulations of $p$, its computational expense grows quickly with the complexity of $\mathcal{Q}$. It is therefore of interest to understand how well simpler variational families can perform. Moving forward, we advocate a workflow in which VI proceeds first with a restricted family $\mathcal{Q}$ (e.g., factorized or location-scale), the accuracy of the inference is checked (here, it would be interesting to develop a symmetry-based check), slight corrections are applied (e.g., Yao:2018; Giordano:2018) and $\mathcal{Q}$ is progressively complexified as necessary.
Acknowledgment
We thank Brian Ward for help with BridgeStan and Isaac Rankin, as well as five anonymous reviewers, for feedback on an early version of this manuscript.